### v1

# The Cross-Lingual Cost: Retrieval Biases in RAG over Arabic-English Corpora

Abstract
Cross-lingual retrieval-augmented generation (RAG) is a critical capability for retrieving and generating answers across languages. Prior work in this context has mostly focused on generation and relied on benchmarks derived from open-domain sources, most notably Wikipedia. In such settings, retrieval challenges often remain hidden due to language imbalances, overlap with pretraining data, and memorized content. To address this gap, we study Arabic-English RAG in a domain-specific setting using benchmarks derived from real-world corporate datasets. Our benchmarks include all combinations of languages for the user query and the supporting document, drawn independently and uniformly at random. This enables a systematic study of multilingual retrieval behavior.
Our findings reveal that retrieval is a critical bottleneck in cross-lingual domain-specific scenarios, with significant performance drops occurring when the user query and supporting document languages differ. A key insight is that these failures stem primarily from the retriever’s difficulty in ranking documents across languages. Finally, we propose a simple retrieval strategy that addresses this source of failure by enforcing equal retrieval from both languages, resulting in substantial improvements in cross-lingual and overall performance. These results highlight meaningful opportunities for improving multilingual retrieval, particularly in practical, real-world RAG applications.
The Cross-Lingual Cost:
Retrieval Biases in RAG over Arabic-English Corpora
Chen Amiraz  Yaroslav Fyodorov  Elad Haramaty  Zohar Karnin  Liane Lewin-Eytan Technology Innovation Institute {chen.amiraz,yaroslav.fyodorov,elad.haramaty,zohar.karnin,liane.lewineytan}@tii.ae
1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as the widely accepted approach for grounding large language models (LLMs) in external knowledge, with most research and development focused on high-resource languages, most notably English. However, many real-world applications, especially in corporate contexts, rely on multilingual corpora, where content spans both high- and low-resource languages. For example, internal knowledge management systems in governmental or legal domains often store content in both a high-resource language like English and the local language, while customer support systems may receive queries in the local language that require retrieving information from a corpus that mixes technical content in both languages. These scenarios introduce cross-lingual complexity, where users interact in a low-resource language while relevant information resides in a corpus containing documents in multiple languages. Prior work has shown that system performance in such cross-lingual settings tends to lag behind monolingual setups, due to challenges across both retrieval and generation Wu et al. (2024); Sharma et al. (2025); Park and Lee (2025). In this work, we focus on the English-Arabic setting – a representative and important case of high- and low-resource language interaction.
Prior work has primarily focused on the generation component Liu et al. (2025); Chirkova et al. (2024), often using multilingual benchmarks derived from Wikipedia, the predominant open-domain source. However, evaluating retrieval in this context poses challenges due to several inherent characteristics: language imbalances, overlap with pretraining data, and the fact that much of Wikipedia’s knowledge is embedded in the model’s parametric memory. In contrast, our work focuses on the less explored retrieval component within a bilingual, domain-specific setting representative of real-world corporate applications.
To this end, we construct benchmarks from UAE corporate datasets with parallel English-Arabic documents. Each benchmark includes a balanced set of English and Arabic queries, with answers grounded in a single language. The languages of the user query and supporting document are selected independently, enabling a systematic analysis of cross-lingual biases. Our findings show that retrieval quality is a key driver of end-to-end performance, with the cross-language setting posing the greatest challenge. Further analysis reveals that the main source of error stems not from the language mismatch between the user query and the supporting document, but from limitations in cross-lingual retrieval within a shared multilingual embedding corpus space.
Finally, we propose a simple mitigation strategy tailored to the identified error source, which retrieves an equal number of passages from each language, resulting in significant improvements in cross-lingual retrieval. The effectiveness of such a basic intervention suggests that there remains substantial room for advancements in this area.
2 Related Work
Cross Lingual Information Retrieval (CLIR) is a critical capability for accessing knowledge across language boundaries, and has gained renewed attention with the rise of cross-lingual retrieval-augmented generation (RAG) systems. These systems typically operate in two phases, retrieval and answer generation. CLIR has historically been done via translation (see Galuščáková et al. (2021) and references within). With the rise of dense retrieval, most leading techniques avoid direct translation and instead embed queries and documents of different languages into the same space Chen et al. (2024); Louis et al. (2025); Wang et al. (2024); Asai et al. (2021b). The improved performance over retrieval tasks was also verified to occur in RAG for question answering by Chirkova et al. (2024) that show an advantage to these direct methods over translation coupled with monolingual retrieval. The different retrieval techniques vary in their training method and data collections, yet all follow the same pattern of embedding the query and the document. They fall into the broader area of Cross Lingual Allignment, where the objective is to align representations of different languages Hämmerl et al. (2024). This broader area and the specifics of the different models are outside the scope of this paper.
For the answer generation phase, the challenge comes from the fact that (1) the user language may not be the same as the retrieved document(s) language, and (2) the documents may be written in multiple languages. Liu et al. (2025) provide a benchmark containing questions that require reasoning. They show that the language difference between the user and document languages can cause issues such as answers in the wrong language. They also show that documents of different languages make cross document reasoning more challenging. Ranaldi et al. (2025) show a simple yet effective method for overcoming both issues; they use a translation service to translate the query and documents to english, then translate the answer. In contrast, Wu et al. (2024) show (on different benchmarks) that this translation-based method breaks down when using lower-quality translation systems, such as medium-scale LLMs. Chirkova et al. (2024) provide practical solutions to the issue of a different user and document language; they highlight comments that when added to the system prompt, result in improved performance. Qi et al. (2025) focus on generation in cross-lingual RAG settings, addressing the influence of retrieved passages both when they are relevant, regardless of their language, or when distracting passages in different languages are provided in the context.
Several studies have examined bias in both retrieval and generation, namely the preference for high-resource languages like English over low-resource ones such as Arabic. Wu et al. (2024) evaluate end-to-end RAG performance across multiple LLMs and show that high-resource languages consistently outperform low-resource ones in both monolingual and cross-lingual settings. They also find that, when relevant documents exist in multiple languages, English passages are more likely to be selected. Sharma et al. (2025) manually constructs a small benchmark over a synthetic corpus to avoid the influence of the parametric memory, and observe a consistent bias favoring the user query language in both stages. Park and Lee (2025) analyze language preferences in both retrieval and generation, highlighting a strong bias toward high-resource languages, especially when the query and document languages match. English is noted as an exception, often outperforming even monolingual configurations – an effect attributed to English dominance in pretraining data.
Most prior work on multilingual RAG, including those cited here, relies on Wikipedia-based datasets and derived benchmarks such as MKQA Longpre et al. (2021), XOR-QA Asai et al. (2021a), and MLQA Lewis et al. (2020). However, Wikipedia introduces several inherent properties: it is significantly richer in English content, has been typically used during the pretraining of both retrievers and generators, and much of its factual knowledge is encoded in the model’s parametric memory. All these factors impact cross-lingual behavior, and in particular, the behavior and influence of retrieval. Chirkova et al. (2024), while focusing on benchmarks derived from Wikipedia, explicitly acknowledge that retrieval performance in multilingual specialized domains remains under-explored.
Thus, our work addresses a gap that has received limited attention by focusing on the retrieval component in a domain-specific, bilingual corporate setting involving a high- and low-resource language pair (English-Arabic). It uses clean multilingual corpora with well-aligned content across both languages, which are unlikely to have been seen during pretraining and represent realistic and practical RAG use cases.
3 Evaluation Pipeline
We use a cross-lingual basic RAG setup focused on English and Arabic. Given a query in either language, its goal is to generate an answer in the same language. The corpus includes documents in both languages, and each query is associated with a ground-truth answer found in one language only. The other language may contain partial or no relevant information.
Our RAG pipeline consists of the standard components: retrieval, re-ranking, and answer generation. Retrieval is performed using dense vector search over a bilingual corpus. We experiment with the multilingual embedding models BAAI BGE-M3111https://huggingface.co/BAAI/bge-m3 (referred to as BGE-M3 from now on) and Multilingual-E5-Large222intfloat/multilingual-e5-large (referred to as M-E5), both of dimension $10241024$, along with the BGE-v2-M3333https://huggingface.co/BAAI/bge-reranker-v2-m3 re-ranker. These models were chosen because they are considered among the top-performing open-source retrievers and re-ranker, and have been widely adopted in previous works. For answer generation, we use Qwen-2.5-14B-Instruct [^4] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct, a large language model with strong multilingual capabilites. During inference, the 20 most relevant passages are retrieved for a given question, then re-ranked based on their relevance and utility for answer generation. The top-5 ranked passages are used to augment the prompt provided to the LLM for answer generation (using Prompt A.3).
3.1 Metrics
An effective RAG system requires success at three stages: retrieving a relevant passage, preserving it through re-ranking, and leveraging it in generation to produce an accurate answer. We analyze the overall end-to-end performance, as well as each component in isolation: retrieval, re-ranking, and generation.
The end-to-end performance and the generation component are evaluated using an answer quality metric, which we refer to as accuracy, based on a semantic equivalence to ground-truth answers provided by our benchmarks (see next section). Specifically, we adopt an LLM-as-a-judge approach to assess correctness, using Claude 3.5 Sonnet to determine whether a generated answer matches the ground-truth reference (see Prompt A.3), following recent work by Zheng et al. (2023). Although LLM-based judgments have faced critique, particularly for relevance assessment Soboroff (2024), prior studies have shown a high correlation with human evaluations in QA contexts. Moreover, the common alternative of strict lexical match is even less reliable in a multilingual setting, as discussed for example in Qi et al. (2025), making a semantic measure more appropriate.
To further support this choice, we validated the metric through human evaluation with native speakers of the tested languages, confirming over 95% agreement between human and automated ratings for both English and Arabic (see Appendix A.1.1 for more details). Given our focus on semantic similarity with respect to the ground truth, we find LLM-as-a-judge to be a practical and reliable measure.
For evaluating the retrieval component, we measure whether the ground-truth answer can be inferred from each retrieved passage. We obtain these relevance judgments using Claude 3.5 Sonnet with Prompt A.3. Based on these relevance labels, we report Hits@20, indicating whether a relevant passage appears among the top $2020$ retrieved results. For reranking, we apply the same procedure and report Hits@5 to measure whether relevant passages appear among the top results of the reranked list. Measuring the presence of relevant passages among the top results is particularly important in a RAG setting, as it reflects whether downstream components have access to the required evidence. The validity of these metrics is supported by their correlation with downstream accuracy, as detailed in Appendix A.1.2.
3.2 Our Benchmarks
We focus on a corporate setting and construct two benchmarks, each based on a separate corpus. Both benchmarks are derived from public websites that contain parallel content in English and Arabic. The first benchmark, Legal, is based on the UAE Legislation website [^5] https://uaelegislation.gov.ae/, which contains 390 laws, with each law described in separate documents in English and Arabic. The second benchmark, Travel, is based on the UAE Ministry of Foreign Affairs website [^6] https://www.mofa.gov.ae/ar-ae/travel-updates, which offers travel-related information for multiple countries, such as visa requirements and embassy contacts. For each country, the information is presented in two parallel documents, one per language.
Having each document available in both languages is essential for our experimental design. In order to build a corpus for each of these two use cases, we assign a document language to each document uniformly at random during corpus construction, ensuring that every document appears in exactly one language within the corpus. The resulting Legal corpus includes roughly 1.5M words, while the Travel corpus contains around 150K words. After building and indexing this bilingual corpus, we proceed to create the benchmark. We used DataMorgana Filice et al. (2025), a synthetic question–answer generation tool, to create query–answer pairs per document, ensuring that each question could be answered using that document alone. The language of each query–answer pair (the user language) is also selected uniformly at random and independently of the document language, resulting in a benchmark that supports systematic evaluation across all language combinations, and allows to identify the source of bias. The final benchmarks include around 1.6K question–answer pairs for Legal and 2K for Travel. Details of the DataMorgana configuration we used to generate our benchmarks are provided in Appendix A.2777The benchmarks will be made publicly available upon publication of this paper..
4 Experiments
We present four experiments, each structured with a description, results, and key conclusions. The first experiment demonstrates that retrieval is a major bottleneck in our bilingual setting. The second identifies the root cause of the retrieval issues as the cross-lingual mismatch between the user language and the document language. The third experiment attributes the performance drop to the presence of a bilingual corpus. Finally, the fourth experiment proposes and evaluates a mitigation strategy to address this issue.
4.1 Retrieval is a Critical Bottleneck
Table 1 presents the results of our first experiment, using the metrics described in Section 3.1. We first measured accuracy without retrieval augmentation for each benchmark. Then, for each of our two embedding models, we evaluated the performance of each system component as well as the overall end-to-end performance.
Specifically, we report Hits@20 for the retrieval phase. For reranking, we report Hits@5 only on examples where retrieval achieved Hits@20 equal to 1, meaning a passage with the answer was passed to the reranker. For generation, we report answer accuracy only on examples where reranking achieved Hits@5 equal to 1, namely where a passage containing the answer was included in the prompt. This analysis helps identify how each phase contributes to the overall end-to-end accuracy.
Benchmark No-RAG Embedder Retrieval Reranking Generation End-to-End
Legal 27±3% BGE-M3 82±2% 89±2% 79±2% 60±2%
M-E5 67±2% 87±2% 79±3% 50±2%
Travel 37±3% BGE-M3 89±1% 97±1% 88±2% 80±2%
M-E5 76±2% 97±1% 85±2% 68±2%
Table 1: No-RAG baseline and RAG component-wise and end-to-end performance. For each benchmark, we first report the baseline answer accuracy using only the user question without retrieval augmentation, referred to as No-RAG. Then, for each embedding model, we report the retriever Hit@20, the reranker Hit@5 conditioned on successful retrievals, the generation answer accuracy conditioned on successful rerankings, and the overall end-to-end answer accuracy. Each value is presented with its 95% confidence interval.
The Legal benchmark represents a domain-specific setting, where questions involve niche topics, so the LLM cannot rely on its parametric memory alone to answer them, as shown by the low accuracy achieved without RAG. This is further confirmed by comparing the end-to-end score in Table 1 with the product of retrieval score, reranking score conditioned on successful retrieval, and generation score conditioned on successful reranking. These values are nearly identical, indicating that the generation phase cannot compensate for failures earlier in the pipeline. The table shows similar results for the Travel benchmark, although the overall accuracy for this case is slightly higher than the product of the component-level conditional scores. This is likely because the Travel corpus includes less specialized knowledge, making it better represented in the LLM’s parametric memory, as also reflected by the performance gap without retrieval.
Looking more closely at the individual components, the reranker performs the best of the three. For both benchmarks with the BGE-M3 embedder, the probability of retrieval failure is comparable to that of generation. With the M-E5 embedder, the retrieval gap is even larger than the generation gap, showing a 12% difference on the Legal benchmark and 9% on Travel. Moreover, for each benchmark, reranking and generation performance are stable across embedders. However, changing retrievers has a significant effect on end-to-end accuracy. These results, taken together, highlight that the retriever is a critical bottleneck and motivate us to focus our efforts on it.
4.2 Cross-Lingual Combinations are the Most Challenging
Benchmark User Lang. Doc Lang. Retrieval Hit@20 End-to-End Accuracy
Legal Arabic Arabic 92±3% 67±5%
Arabic English 89±3% 68±4%
English Arabic 56±5% 30±5%
English English 86±3% 68±4%
[…]
Cross-lang. 72±3% 49±3%
Travel Arabic Arabic 93±2% 85±3%
Arabic English 91±2% 78±4%
English Arabic 79±4% 72±4%
English English 93±2% 84±3%
[…]
Cross-lang. 85±2% 75±3%
(a) BGE-M3 embedder
Benchmark User Lang. Doc Lang. Retrieval Hit@20 End-to-End Accuracy
Legal Arabic Arabic 88±4% 67±5%
Arabic English 51±4% 38±4%
English Arabic 41±5% 22±4%
English English 88±3% 70±4%
[…]
Cross-lang. 46±3% 30±3%
Travel Arabic Arabic 90±3% 86±3%
Arabic English 55±4% 38±4%
English Arabic 64±4% 62±4%
English English 95±2% 86±3%
[…]
Cross-lang. 59±3% 50±3%
(b) M-E5 embedder
Table 2: Performance across language combinations. Results are presented for each embedder, benchmark and for each of the four possible user–document language combinations. In addition, we report same-language and cross-language scores, defined as the mean scores over combinations where the user and document languages match or differ, respectively.
Next, we compare the retrieval and end-to-end performances on each of the four user-document language combinations. The results for the BGE-M3 and M-E5 embedders are presented in Tables 2 and 2, respectively.
The tables reveal that cross-lingual scenarios, where the user query and the supporting document are in different languages, consistently underperform compared to same-language settings in both retrieval and end-to-end performance. For the BGE-M3 embedder, a significant decline in retrieval performance is observed only when the user language is English and the document language is Arabic, with drops of 33% in the Legal benchmark and 14% in Travel compared to the same-language configuration. A similar pattern appears in the final accuracy, with decreases of 38% and 12%, respectively. Notably, the reverse cross-lingual setting does not exhibit any statistically significant degradation for BGE-M3.
In contrast, the M-E5 embedding exhibits an even larger performance drop across both cross-lingual settings. Specifically, retrieval Hit@20 decreases by 42% on the Legal benchmark and by 33% on Travel, compared to their same-language counterparts. These retrieval declines also propagate to the end-to-end accuracy, resulting in drops of 38% for Legal and 36% for Travel.
In what follows we dive deeper to discover the cause behind this gap.
4.3 The Source of the Cross-Lingual Failure
Notice that in our current setup, referred to from now on as the direct setting, multilinguality arises from both the corpus and the queries. This creates two main challenges. First, the retriever must handle cross-lingual queries, ranking documents in language X given a query in language Y. Second, it must rank documents across languages without introducing bias toward either a high-resource language or the user language.
To determine which of these challenges is primarily responsible for the observed failures, we conducted the following experiment. We modified the retriever from the direct setting to search only within the correct language. Specifically, for a document language X, the language-oracle retriever returns the top results exclusively in language X, completely excluding language Y. Hence, while the user query can still be in a different language, all retrieved passages, and therefore all subsequent stages of re-ranking and generation, operate exclusively on monolingual documents. This setup allows us to isolate the challenges of multilingual query embeddings while avoiding issues related to multilingual document embeddings. We stress that the oracle is used only for analysis purposes, since in practice we do not have access to the document language ahead of time. The first two bars in each subfigure of Figure 1 present the Hit@20 performance of the direct and language-oracle retrievers, broken down by user-document language combinations, as well as overall.
(a) Legal benchmark – BGE-M3 embedder
(b) Legal benchmark – M-E5 embedder
(c) Travel benchmark – BGE-M3 embedder
(d) Travel benchmark – M-E5 embedder
Figure 1: Retrieval Hit@20 scores across benchmarks and embedders. Each figure corresponds to a specific combination of benchmark and embedding. Bars represent retrieval Hit@20 scores in percentages, with 95% confidence intervals shown as black error lines. Different retrieval policies are distinguished by color and texture. Results are grouped by benchmark segments defined by the user-document language combination, as well as the overall benchmark retrieval performance.
We observe two clear phenomena. First, the language-oracle retriever achieves nearly identical performance across all user and document language pairs, suggesting there are essentially no failures related to the query embeddings. In contrast, the gap between the direct and language-oracle retrievers can be substantial in many cross-lingual cases. This indicates that the main source of failure lies in the multilingual document embeddings, specifically the retriever’s ability to rank documents across languages.
The results suggests that while semantic similarity is well captured within a single language, the embeddings struggle in cross-lingual settings. For instance, BGE-M3 appears to favor English passages when the user query is in English, while M-E5 may exhibit a tendency to prefer passages in the same language as the user query.
4.4 Mitigating Cross-lingual Failings
These results raise an important question: can multilingual retrievers be used reliably on mixed-language corpora without further tuning? To explore this question, we introduce the balanced retriever, which enforces an equal selection of documents from each language by retrieving 10 passages in Arabic and 10 in English. We evaluate this approach in the same experimental setup described earlier. The third and final bar in each subfigure of Figure 1 presents its results. Notice that while the language-oracle retriever is infeasible, it serves as an upper bound for what the balanced retrieval could achieve. In practice, we observe that the balanced retriever does not significantly degrade performance on same-language cases while providing substantial improvements in cross-lingual cases. Notably, the balanced retriever yields more consistent results across the different combinations of user and document languages, unlike the direct setting, which favors the same-language combinations at the expense of cross-language ones. Moreover, this strategy leads to a significant improvement in overall retrieval accuracy across benchmarks and embedders, with consistent gains of around 3-5% for BGE-M3 and approximately 20% for M-E5.
These findings suggest that even simple strategies can help mitigate retrieval biases, indicating that debiasing is feasible despite inherent biases in the embedding model.
5 Conclusions
This work highlights retrieval as a critical bottleneck in multilingual RAG systems applied to domain-specific corpora. While prior studies have identified and focused on generation as the main limitation in cross-lingual RAG, their conclusions are primarily based on Wikipedia-derived benchmarks. Since multilingual retrievers such as BGE-M3 and multilingual-E5-large are trained on similar open-domain data, they exhibit strong performance in those settings. In contrast, our domain-specific benchmarks expose substantial retrieval weaknesses that remain obscured in such evaluations, underscoring the need to revisit cross-lingual retrieval in practical, real-world RAG scenarios.
Our analysis shows that performance degrades most in cross-lingual settings where the user and document languages differ, with drops of 30–50% compared to same-language configurations. Using an oracle retriever restricted to the correct language, we isolate the primary source of failure as the retriever’s difficulty in ranking documents across languages. That is, while the retriever performs well within a single language, it struggles when comparing passages across languages, often favoring those in the query’s language. We further observe that different embedders exhibit weaknesses in different cross-lingual settings. This highlights the potential to improve training by explicitly targeting cross-lingual robustness and narrowing the gap with same-language performance.
Lastly, we demonstrate that a simple mitigation of retrieving a balanced number of documents per language, can significantly improve cross-lingual performance and even improve overall results. This finding points to meaningful opportunities for mitigating multilingual retrieval biases, particularly in real-world applications. However, applying such a balancing approach in practical settings with non-uniform language distributions or more than two languages remains an open challenge and requires further investigation.
Acknowledgments
We thank our colleagues at AI71, and in particular Abdelrahman Ibrahim, Amr Ali Abugreedah, Mohamad Salah, Saleem Hamo, Kirollos Sorour, Imran Moqbel, and Anas AlHelali, whose native proficiency in Arabic was instrumental in the annotation process used to validate the evaluation metrics in our pipeline.
Appendix A Appendix
A.1 Metric Evaluation
A.1.1 Answer Accuracy
To validate our answer accuracy metric across languages, we performed the following procedure. First, we used samples from English and Arabic editions of Wikipedia to construct two benchmarks of 100 examples each, using DataMorgana Filice et al. (2025). We then applied a standard RAG pipeline to generate answers using Falcon-3-10B. The generated answers were then compared to reference answers using our LLM-as-a-judge-based accuracy metric, as described in the main text. This setting was intentionally selected to produce a mix of correct and incorrect answers, ensuring a meaningful evaluation of the metric.
[…]
A.1.2 Retrieval Hit@20
Now that we trusted our LLM-based accuracy metric, we moved to validating whether our Hits@20 metric, which also uses LLM judgments, effectively captures success in the retrieval step. Towards this goal, we analyzed the downstream accuracy as a function of the Hit@20 score. This analysis focused on the Legal benchmark, where the no-RAG accuracy is relatively low (27%), making it easier to observe the impact of retrieval quality. Table 3 reports these results for the BGE-M3 and M-E5 embedders.
End-to-End Accuracy
Retrieval Hit@20 BGE-M3 M-E5
0 10±3% 9±2%
1 79±2% 79±3%
Overall 60±2% 50±2%
Table 3: End-to-End Accuracy as Function of Our LLM-based Hit@20. Each cell shows the average accuracy along with its 95% confidence interval. Columns correspond to retrieval embedders; rows indicate evaluation segments: instances with Hit@20 = 0, Hit@20 = 1, and overall accuracy.
As shown in Table 3, downstream accuracy was indeed low when the Hits@20 metric indicates failure, confirming that our LLM-based Hits@20 reliably identifies cases where retrieval has failed. Specifically, accuracy dropped to approximately 9% when no relevant passage was identified by the metric, which is considerably lower than the 27% accuracy observed without retrieval augmentation. Furthermore, we observed consistent patterns across retrievers: although the BGE-M3 retriever differed markedly in overall quality from the M-E5 retriever, their downstream accuracy as a function of retrieval quality showed only minor differences, likely attributable to statistical noise. These findings validate our Hits@20 metric as a reliable measure of retrieval effectiveness, demonstrating that higher scores are strongly associated with improved downstream accuracy.
A.2 DataMorgana Configuration
The following describes the configuration used to construct both the Legal and Travel benchmarks. In both cases, DataMorgana was configured in non-conversational mode, supporting single-turn question answering only.
DataMorgana allows the definition of multiple parallel question categorizations, each selected independently of the rest of the configuration, including other categories and the document language. The question categorizations were defined as follows:
  • Language: The user language was set to Arabic in 50% of the cases and English in 50%.
  • Formulation: The question was phrased as:
    • Concise natural language: 40% of cases.
    • Verbose natural language: 20% of cases.
    • Short search query: 25% of cases.
    • Long search query: 15% of cases.
  • Linguistic similarity: In 50% of the cases, the phrasing was similar to that found in the corpus, and in the remaining 50%, it had a greater linguistic distance.
  • Question type: Questions were evenly split between factoid (50%) and open-ended (50%).
  • User need:
    • For the Legal benchmark, 50% of the questions simulated a user seeking specific legal advice, while the other 50% simulated a user asking out of general curiosity.
    • For the Travel benchmark, the user type was distributed as follows: UAE user in 20% of the cases, Non-UAE user in an additional 30%, and Undisclosed citizenship in the remaining 50%.
A.3 Prompts
In this section, we provide all the prompts used in our experiments. Prompt A.3 was used for answer generation. It is based on the guidelines proposed by Chirkova et al. (2024) for prompting RAG systems in multilingual scenarios. Prompt A.3 was used to evaluate the accuracy of the generated answer. Prompt A.3 was used to evaluate retrieval Hit@20 and reranking Hit@5.
Prompt A.1: RAG generation
System.
Answer the question based on the given passages below. Elaborate when answering, and if applicable provide additional helpful information from the passages and only from the passages. Do not refer to the passages, just state the information. You MUST answer in the SAME LANGUAGE as the QUESTION LANGUAGE, regardless of the language of the passages. Answering in the same language as the user is asking their question is crucial to your success. If the question is in English, the answer must also be in English. If the question is in Arabic, the answer must also be in Arabic. Write all named entities in the same language and same alphabet as the question language.
User.
# Passages: passage 1: <Passage 1> passage 2: <Passage 2> passage 3: <Passage 3> # Question: <Question>
Prompt A.2: Generated answer evaluation Based on the question and the golden answer, judge whether the predicted answer has the same meaning as the golden answer. Return your answer in the following format: <same_meaning>True/False</same_meaning>. <question> … </question> <golden_answer> … </golden_answer> <predicted_answer> … </predicted_answer>
Prompt A.3: Retrieval evaluation You are given a **question**, a **ground truth answer**, and a list of **passages**. Your task is to return the **list of passage indices** that can directly answer the question **by containing the ground truth answer** (i.e., the passage includes a perfect match to the information expressed in the ground truth). Please follow these rules: - A passage should be included only if it **clearly expresses or contains the ground truth answer**. - Do **not include passages** that are only loosely related or provide background information. - Your response **must be valid Python list syntax**, e.g., [3, 5, 9]. - Do **not add any explanation** outside the list. **Question**: <Question> **Ground Truth Answer**: <Answer> **Passages**: Passage 1: <Passage 1 content> Passage 2: <Passage 2 content> Passage 3: <Passage 3 content>